A photograph machine is capable of making 50 copies each minute.
(a) If 4 identical copiers are used simultaneously how long would it take to make a total of
50 copies?
(b) How many copiers would be needed to make 6000 copies in 15 minutes?
1
Expert's answer
2012-06-22T09:24:17-0400
(a) If 4 identical copiers are used simultaneously how long would it
take to
make a total of
50 copies?
t = ] N / (n * C) [ = ] 50 / (4 * 50) [ = 0.25
min = 15 sec.
(b) How many copiers would be needed to make 6000 copies in 15
minutes?
n = N / (C * t) = 6000 / (50 * 15) = 8.